# ICONS: Influence Consensus for Vision-Language Data Selection

### v2

@@ 2.4 Generalist: Cross-Task Influence Consensus
Our consensus-based voting strategy identifies training samples that consistently show a high influence score across various tasks. For each task $k$, we set a task-specific threshold $\tau_{k}$ at the $((1-p)\times 100)$-th percentile of the influence score distribution within that task, where $p$ represents the proportion of top samples to consider ($p=0.2$ in our main experiments). A sample ${\bm{z}}_{i}$ is considered important for task $k$ if its influence score $I_{k}({\bm{z}}_{i})\geq\tau_{k}$. Formally, we define the voting-based influence score for sample $z_{i}$ on task $k$ as:
$I_{\text{vote}}({\bm{z}}_{i})=\sum_{k=1}^{K}\mathbbm{1}[I_{k}({\bm{z}}_{i})\geq\tau_{k}],$ (4)
where $\mathbbm{1}[\cdot]$ denotes the indicator function, and $I_{k}({\bm{z}}_{i})$ represents the influence score of sample ${\bm{z}}_{i}$ for task $k$. The final score for each sample is the total number of votes it receives across all tasks, ranging from 0 (indicating no importance for any task) to $K$ (indicating consistent importance for all tasks). Influence Consensus identifies a unique subset of samples that are consistently influential across multiple tasks, rather than those that might be highly influential in only a few tasks. After computing the aggregated influence scores, we rank the training samples and select the top $p\%$ as final subset for model training.
While directly combining validation data from all target tasks and computing a single influence matrix (Direct-merge Selection, §3.5) might seem intuitive, we adopt this consensus-based approach for three reasons. First, specialist selection provides a practical performance upper bound, as the performance achieved by selecting the top $p$of data specifically for a single task represents the theoretical maximum achievable with that data budget, serving as a critical baseline for evaluating generalist selection trade-offs (§3.5). Second, this approach is scalable and naturally supports continual learning scenarios, as influence matrices for new tasks or benchmarks can be computed independently without recomputing training gradients, thereby reducing computational costs. Third, computing influence matrices independently for different tasks give us better understanding of cross-task influence dynamics. Our analysis (§3.3) reveals that the influence ranking of training data vary significantly across different tasks, with highly influential samples for one task often showing limited impact on others. This further provides insights to design more effective consensus mechanisms for generalist selection while maintaining computational tractability.
@@
CLIP-Score [35] 73.4 51.4 43.0 65.0 54.7 85.3 1331.6 55.2 52.0 66.2 91.2
EL2N [33] 76.2 58.7 43.7 65.5 53.0 84.3 1439.5 53.2 47.4 64.9 92.0
Perplexity [28] 75.8 57.0 47.8 65.1 52.8 82.6 1341.4 52.0 45.8 68.3 91.6
SemDeDup [1] 74.2 54.5 46.9 65.8 55.5 84.7 1376.9 52.2 48.5 70.0 92.6
D2-Pruning [26] 73.0 58.4 41.9 69.3 51.8 85.7 1391.2 65.7 57.6 63.9 94.8
Self-Sup [38] 74.9 59.5 46.0 67.8 49.3 83.5 1335.9 61.4 53.8 63.3 93.4
Self-Filter [5] 73.7 58.3 53.2 61.4 52.9 83.8 1306.2 48.8 45.3 64.9 90.9
COINCIDE [18] 76.5 59.8 46.8 69.2 55.6 86.1 1495.6 63.1 54.5 67.3 97.4
@@
Table 1: Baseline Comparisons. Performance comparison of different data selection approaches when trained on 20% of the LLaVA-665K dataset. Models are evaluated across diverse multimodal benchmarks, with tasks ranging from visual question answering (VQAv2, GQA) to multilingual understanding (MMBench). The best and second best results for each benchmark are shown in bold and underlined, respectively. Our method ICONS achieves the highest overall relative performance (98.6%), consistently outperforming existing approaches including COINCIDE (97.4%) and D2-Pruning (94.8%), while methods like EL2N, Perplexity, and CLIP-Score show limited effectiveness with relative performance around 91-92%.
3 Experiments
In this section, we first discuss our experiment setup and evaluation benchmarks (§3.1). Then we compare it with the latest state-of-the-art methods (§3.2). Next, we provide detailed analysis of our method’s behavior and effectiveness across different dimensions (§3.3). We further evaluate the transferability of our selection approach across both unseen tasks and model architectures (§3.4). Finally, we conduct ablation studies to understand the contribution (§3.5).
3.1 Evaluation Test-Bed
Dataset & Model.  We evaluate ICONS on visual instruction tuning dataset LLaVA-665K [22]. We use LLaVA-v1.5-7b-lora model [22] with a default size of 7B parameters unless otherwise specified. In all experiments, we train the models for one epoch following the official finetuning hyperparameters, which reduces computational overhead while maintaining selection quality.
Target Tasks.  We evaluate the effectiveness of ICONS with a wide range of multimodal benchmarks (Tab. 2) that test different capabilities of vision-language models. The benchmarks include: 1) Multiple-choice understanding: MMBench [47] and MME [7], which assess the model’s ability to select correct answers from given options; 2) Visual question answering: VQAv2 [9], GQA [13], and VizWiz [10], which test basic visual reasoning; 3) Text understanding in images: TextVQA [37], which evaluates OCR capabilities; 4) Scientific reasoning: ScienceQA [25], which tests knowledge-grounded question answering; 5) Open-ended generation: LLaVA-W Bench [23], which evaluate free-form response generation; 6) Factual consistency: POPE [21], which measures hallucination.
MME POPE SQA-I MMBench
en cn
$|\mathcal{D}_{\text{val}}|$ 986 500 424 1,164 1,164
$|\mathcal{D}_{\text{test}}|$ 2,374 8,910 4,241 1,784 1,784
Task Type Y/N Y/N MCQ MCQ MCQ
VQAv2 GQA VizWiz TextVQA LLaVA-W
$|\mathcal{D}_{\text{val}}|$ 1,000 398 8,000 84 84
$|\mathcal{D}_{\text{test}}|$ 36,807 12,578 8,000 5,000 84
Task Type VQA VQA VQA VQA VQA
3.2 Comparisons with Baselines
We compare our ICONS against several baseline approaches, including random selection. We evaluate CLIP-Score [35] for measuring image-text alignment, EL2N [33] based on embedding L2 norms, and Perplexity [28] using language model scores. We also compare against SemDeDup [1] for semantic deduplication and D2-Pruning [26] for distribution-aware pruning. Additional baselines include Self-Sup [38] leveraging self-supervised signals, while Self-Filter [5] and COINCIDE [18] are the approaches specifically designed for vision-language data selection. All methods are evaluated using a 20% sampling ratio of the original dataset, with results presented in Tab. 1. We reference the baseline results from COINCIDE [18].
@@ 3.2 Comparisons with Baselines
We compare our ICONS against existing data selection approaches in Tab. 1. Our method achieves the best overall performance with 98.6% Rel., outperforming all baselines with LLaVA-ICONS-133K, 20% of the training data. Remarkably, we achieve comparable or better performance than training with the entire LLaVA-665K on several tasks: SQA-I (70.8 vs. 68.4), MME (1485.7 vs. 1476.9), VizWiz (50.1 vs. 47.8) and POPE (87.5 vs. 86.4). Among the baselines, COINCIDE achieves strong performance (97.4% Rel.), though it falls short of ICONS on key tasks. Other methods like EL2N, Perplexity, SemDeDup are less effective, with relative scores around 91-92%, showing limitations in preserving performance.
Task Full Specialist Generalist Delta (%)
VQAv2 [9] 79.1 77.1 76.3 1.04
GQA [13] 63.0 61.1 60.7 0.65
VizWiz [10] 47.8 53.1 50.1 5.65
SQA-I [25] 68.4 69.8 70.8 -1.43
TextVQA [37] 58.2 55.7 55.6 0.18
POPE [21] 86.4 86.6 87.5 -1.04
MME [7] 1476.9 1506.1 1485.7 1.35
MMBench (en) [47] 66.1 66.0 63.1 4.34
MMBench (cn) [47] 58.9 56.4 55.8 1.06
LLaVA-W Bench [23] 67.9 67.1 66.1 1.49
Table 3: Single-task Selection (Specialist) vs. Consensus-aware Multi-task Selection (Generalist). The single-task data selection approach selects 20% of data per task, while our consensus-aware multi-task data selection approach selects a total of 20% data across all tasks. The Delta (%) column represents the relative drop in performance from specialist to generalist.

### v3

@@ 2.4 Generalist: cross-task influence consensus
Our goal is to identify a training set subset $\mathcal{S}\subset\mathcal{D}$ of size $M\ll N$ such that its performance across all tasks remains high as defined by Eqn. 1. There are multiple ways to tackle this, depending on the assumptions one makes about the task-specific influence scores $\bar{I}_{k}({\bm{z}}_{i})$. The simplest approach is to merge together all the different tasks’ validation sets $\mathcal{D}_{\text{val}}^{k}$ (normalizing for their different sizes) and compute the total influence score for a training example ${\bm{z}}_{i}$ as:
$I_{\text{Merge}}({\bm{z}}_{i})=\sum_{k=1}^{K}\bar{I}_{k}({\bm{z}}_{i}).$ (3)
A similar aggregation approach is the one suggested in LESS xia2024less :
$I_{\text{Max}}({\bm{z}}_{i})=\max_{k=1,\dots K}\bar{I}_{k}({\bm{z}}_{i}),$ (4)
i.e., the influence of the data is measured as its highest influence on any tasks. The set of $M$ training examples with the highest aggregated influence scores would be selected for inclusion in the training set $\mathcal{S}_{\text{merge}}$ (correspondingly, $\mathcal{S}_{\text{max}}$). Both approaches, however, require that the influence scores $\bar{I}_{k}$ be well-calibrated across the different tasks; as we show in §3.3 this may not necessarily be the case.
An alternative approach which does not require directly comparing influence scores $\bar{I}_{k}$ across tasks $k$ is to leverage the relative rank of the training examples within each task. Concretely, we compute $\text{rank}_{k}({\bm{z}}_{i})$ for each example ${\bm{z}}_{i}$ relative to other examples for task $k$ according to their influence scores (higher influence scores correspond to lower rank). We can have a couple of options. First, we can select the training subset either using the Round Robin (RR) approach ivison2025large where we iterate over tasks and select the lowest-rank example which has not yet been selected to add to our set $\mathcal{S}_{\text{RR}}$. Alternatively, we can select the training subset $\mathcal{S}_{\text{MinRank}}$ such that all the examples within it have a low rank for some task $k$. Mathematically, albeit somewhat confusingly, this corresponds to:
$\mathcal{S}_{\text{MinRank}}=\operatorname*{arg\,max}_{\mathcal{S}\subset\mathcal{D},|\mathcal{S}|=M}\min_{\begin{subarray}{c}\text{task }k\\ \text{example }{\bm{z}}_{i}\notin S\end{subarray}}\text{rank}_{k}({\bm{z}}_{i}),$ (5)
i.e., all examples that are not included in $\mathcal{S}$ would have high relative ranks $\text{rank}_{k}({\bm{z}}_{i})$ for all tasks $k$. However, this approach does not consider the potential interplay between tasks. Recall that in Eqn. 1 we aim to maximize sum of the relative performance across all tasks $k$; thus, if a training example is beneficial for multiple tasks, we may want to include it even at the expense of a lower-ranked example for a different task $k$. Thus, we introduce a simple consensus-based voting strategy that identifies training samples that consistently show a high influence score across various tasks. Concretely, we leverage the specialist training sets $\mathcal{S}_{k}$ as defined in §2.3 consisting of the $M$ highest-influence training examples for each task. We then select a combined training set as follows:
$S_{\text{ICONS}}=\operatorname*{arg\,max}_{\mathcal{S}\subset\mathcal{D},|\mathcal{S}|=M}\sum_{k=1}^{K}\sum_{{\bm{z}}_{i}\in S}\mathbbm{1}[{\bm{z}}_{i}\in S_{k}]$ (6)
This simple approach offers a key advantage: it does not rely on calibration of influence scores across tasks, and does not make any a-priori assumptions about the relationship between tasks (e.g., that every task needs to have its highest-scoring examples included in the combined training set). Our generalist stage converts each task’s ranked list into a binary vote (“above threshold” or not) and then combines these votes across tasks, eliminating the need for task-specific normalization. As a result, the selection remains insensitive to scale differences while still capturing relative importance within each task. Meanwhile, a training sample is selected only when several tasks independently rank it as influential, preventing over-representation of single-task outliers and ensuring the cross-task utility.
3 Experiments
In this section, we first discuss our experiment setup and evaluation benchmarks (§3.1). We then present our main results by comparing ICONS with the state-of-the-art methods (§3.2), followed by analysis of different selection strategies (§3.3). We further evaluate the transferability of ICONS3.4). Lastly, we provide analyze performance trends under different selection ratios (§3.5).
3.1 Evaluation test-bed
Datasets & model.  We apply ICONS on major visual instruction tuning (VIT) training datasets: LLaVA-665K liu2024improved , Cambrian-7M tong2024cambrian and Vision-Flan-186K visionflan . The majority of our analysis and ablation experiments are conducted on LLaVA-665K. For our experiments, we use the LLaVA-v1.5 model (liu2024improved, ) checkpoint after Stage 1 (pre-training for feature alignment) as defined in the original LLaVA training pipeline, with a default size of 7B parameters and LLaVA-665K unless otherwise specified. This checkpoint333 llava-v1.5-mlp2x-336px-pretrain-vicuna-7b-v1.5, which has no prior exposure to the visual instruction tuning data. corresponds to the model after training the projector but before any visual instruction tuning in Stage 2. Importantly, this model has not been exposed to the LLaVA-665K VIT dataset prior to the data selection process. In all experiments, we train the models for one epoch following the official finetuning hyperparameters using LoRA. More details on computation, including hardware specifications and runtime are in Appendix A.
Target tasks.  We evaluate ICONS across diverse multimodal benchmarks (Appendix C, Tab. 5) that test different capabilities of vision-language models: 1) Multiple-choice understanding: MMBench (mmbench, ) and MME (mme, )  [^4] For MME, we focus on its perception section following coincide , which evaluates vision capablities. 2) Visual question answering: VQAv2 (vqav2, ), GQA (gqa, ), and VizWiz (vizwiz, ); 3) Text understanding in images: TextVQA (textvqa, ); 4) Scientific reasoning: ScienceQA (sqa, ); 5) Open-ended generation: LLaVA-W Bench (llava, ); 6) Factual consistency: POPE (pope, ).
Baselines.  We compare our ICONS against several baselines, including random selection, CLIP-Score clip for measuring image-text alignment, EL2N el2n based on embedding L2 norms, and Perplexity perplexity using language model scores. We also compare against SemDeDup semdedup for semantic deduplication and D2-Pruning (d2, ) for distribution-aware pruning. Additional baselines include Self-Sup (selfsup, ) leveraging self-supervised signals, while Self-Filter (self-filter, ) and COINCIDE (coincide, ) are designed for vision-language data selection. We reference LLaVA-665K baseline results from COINCIDE coincide . Additionally, we compare with representation-based data selection baseline (RDS) ivison2025large ; xia2024less .
@@ 3.1 Evaluation test-bed
CLIP-Score clip 73.4 51.4 43.0 65.0 54.7 85.3 1331.6 55.2 52.0 66.2 91.2
EL2N el2n 76.2 58.7 43.7 65.5 53.0 84.3 1439.5 53.2 47.4 64.9 92.0
Perplexity perplexity 75.8 57.0 47.8 65.1 52.8 82.6 1341.4 52.0 45.8 68.3 91.6
SemDeDup semdedup 74.2 54.5 46.9 65.8 55.5 84.7 1376.9 52.2 48.5 70.0 92.6
D2-Pruning d2 73.0 58.4 41.9 69.3 51.8 85.7 1391.2 65.7 57.6 63.9 94.8
Self-Sup selfsup 74.9 59.5 46.0 67.8 49.3 83.5 1335.9 61.4 53.8 63.3 93.4
Self-Filter self-filter 73.7 58.3 53.2 61.4 52.9 83.8 1306.2 48.8 45.3 64.9 90.9
COINCIDE coincide 76.5 59.8 46.8 69.2 55.6 86.1 1495.6 63.1 54.5 67.3 97.4
RDS ivison2025large ; xia2024less 75.1 57.9 48.6 68.0 54.9 86.3 1393.8 61.2 52.7 63.7 95.2
@@
Table 1: Selection results on LLaVA-665K. Performance comparison of different data selection approaches when trained on 20% of the LLaVA-665K dataset. The best and second best results for each benchmark are shown in bold and underlined, respectively. Our method ICONS achieves the highest overall Rel. (98.6%), consistently outperforming existing approaches including COINCIDE coincide (97.4%) and D2-Pruning d2 (94.8%).
3.2 Main results
@@
LLaVA-665K selection.  As shown in Tab. 1, ICONS achieves the best overall performance with 98.6% Rel. on LLaVA-665K, outperforming all baselines with LLaVA-ICONS-133K, 20% of the training data. Remarkably, we achieve comparable or better performance than full dataset training on several tasks: SQA-I (70.8 vs. 68.4), MME (1485.7 vs. 1476.9) and POPE (87.5 vs. 86.4). While COINCIDE achieves strong performance (97.4% Rel.), it falls short of ICONS on key tasks. Approaches like EL2N, Perplexity, SemDeDup achieve only 91-92% Rel., showing limitations in preserving performance.
Dataset #Data Method VQAv2 GQA VizWiz SQA-I TextVQA POPE MME MMBench LLaVA-W Rel. (%)
en cn Bench
Vision-Flan-186K 186k Full 68.0 49.2 41.7 60.8 50.4 83.4 1,263.2 52.6 45.9 63.3 100.0
37k Random 64.1 45.8 37.5 58.7 45.3 82.9 1,079.8 46.5 39.6 58.7 91.6
ICONS (ours) 67.4 48.8 44.1 60.2 49.9 83.0 1,252.5 51.9 45.4 63.1 99.8
Cambrian-7M 7,068k Full 80.2 62.9 58.4 75.3 60.9 86.5 1,524.6 69.1 58.9 67.6 100.0
1,414k Random 74.2 57.5 61.9 71.0 57.1 86.4 1,465.7 63.3 49.6 70.4 95.4
ICONS (ours) 79.6 62.1 60.7 73.9 59.8 86.2 1,503.1 67.8 55.8 67.0 98.8
Table 2: Selection results on Vision-Flan-186K and Cambrian-7M. Performance comparison of different data selection approaches when trained on 20% of the Vision-Flan-186K visionflan and Cambrian-7M tong2024cambrian datasets. ICONS achieves strong performance (99.8% and 98.8% Rel. respectively) while using only 20% of the training data, significantly outperforming random selection which is one of the strongest baselines, and approaching full performance.
Cambrian-7M & Vision-Flan-186K selection.  We further provide results on Vision-Flan-186K and Cambrian-7M in Tab. 2. On Vision-Flan-186K, our method achieves near-full performance (99.8% Rel.) using just 37k samples, significantly outperforming random selection (91.6%) across tasks. Similarly, on Cambrian-7M, ICONS maintains a strong performance (98.8% Rel.) with 1,414k samples, while random selection achieves 95.4%. These results demonstrate that our approach scales effectively to both small and large datasets, consistently preserving model capabilities while drastically reducing the training data required.
Comparisons with representation-based data selection.  We compare our method against RDS (Representation-based Data Selection)ivison2025large ; xia2024less , a strong baseline in language-only instruction tuning. RDS computes training-validation similarity using final-layer representation of the last token in each sequence instead of gradients. For a fair comparison, we use the same influence matrix formulation (Eqn.2) and apply majority voting to reach influence consensus. Our method consistently outperforms RDS across all tasks, particularly those requiring perceptual grounding – e.g., higher scores on GQA (60.7 vs. 57.9), SQA-I (70.8 vs. 68.0), and MME (1485.7 vs. 1393.8). While RDS is effective in selecting large-scale text-only data (e.g. TULU-2/3 ivison2023camels ; lambert2024t ), its evaluation has largely focused on language-only tasks, where semantic similarity alone is often sufficient. In contrast, vision-language tasks demand alignment between modalities, where representation similarity is limited as it only reflects current embedding proximity and gradient-based approaches directly estimate each sample’s contribution to the validation loss. Our gradient-based approach directly estimates each sample’s impact on validation loss, capturing cross-modal training dynamics and prioritizing impactful training points. We further provide qualitative comparisons in Appendix §I.
3.3 Analysis of aggregation strategies
Ablations.  As introduced in §2.4, we explore multiple strategies for aggregating task-specific influence rankings into a single compact subset. We compare our majority voting approach (Vote) with different aggregation approaches for combining task-specific influence scores: 1) score-based methods (Merge, Max) and their normalized variants (Merge-GausNorm, Merge-SumNorm), 2) rank-based methods (Round Robin, MinRank). Our voting-based strategy outperforms all alternatives (Tab. 3), achieving the highest overall Rel. (98.6%). This shows that building a cross-task consensus via majority voting is a simple yet effective strategy for identifying consistently influential examples across tasks, without assuming calibration or comparability of scores.
Aggregation VQAv2 GQA VizWiz SQA-I TextVQA POPE MME MMB (en) MMB (cn) LLaVA-W Rel. (%)
Full 79.1 63.0 47.8 68.4 58.2 86.4 1476.9 66.1 58.9 67.9 100
Merge 75.7 59.6 47.9 65.5 55.5 86.0 1422.1 59.0 51.0 66.2 96.4
Max 75.2 59.8 48.1 66.2 55.5 85.5 1470.7 58.3 51.8 66.2 96.1
Merge-GausNorm 75.1 60.1 46.4 69.8 54.5 85.6 1482.6 58.9 52.5 66.3 96.8
Merge-SumNorm 75.5 59.1 51.7 68.7 43.5 87.1 1478.3 59.5 50.9 69.8 95.3
Round Robin 75.4 59.1 48.3 70.6 55.2 86.6 1474.5 61.6 51.5 66.9 96.7
MinRank 75.2 59.0 49.7 70.4 55.1 86.9 1456.3 61.1 52.4 68.4 97.1
Vote (ours) 76.3 60.7 50.1 70.8 55.6 87.5 1485.7 63.1 55.8 66.1 98.6
Table 3: Comparison of aggregation approaches. Performance of different influence aggregation methods when selecting 20% of the LLAVA-665K dataset. Our proposed aggregation approach (Vote) consistently achieves the best overall performance (98.6% Rel.), outperforming both score-based (Merge, Max), their noramlized variants (Merge-GausNorm, Merge-SumNorm) and rank-based (Round Robin, MinRank) baselines.
Limitations of score-based aggregation.  Score-based methods like Merge (Eqn. 3) and Max (Eqn. 4) assume calibrated influence scores across tasks, which is rarely the case. We observe substantial variation in the distribution of influence scores across tasks with standard deviations spanning from $8.15\times 10^{-3}$ (MME) to $1.26\times 10^{-3}$ (VQAv2), indicating that influence scores for MME are much more spread out, while those for VQAv2 are tightly concentrated. Similarly, mean influence scores vary in both magnitude and sign: MME has a relatively high positive mean ($1.68\times 10^{-3}$), while tasks like POPE ($-2.83\times 10^{-4}$) and GQA ($-8.89\times 10^{-5}$) have negative means. These divergent patterns show that certain tasks have much wider influence scores distributions, making a sample helpful for one task but neutral or harmful for another. Aggregating raw scores thus biases selection toward tasks with higher variance or skewed means. To address this calibration issue, we experimented with normalization strategies: Merge-SumNorm, which rescales each task’s influence scores by dividing them by a task-specific normalization factor (i.e., sum), and Merge-GausNorm, which normalizes the scores using task-wise mean and standard deviation before averaging:
$I_{\text{Merge-SumNorm}}({\bm{z}}_{i})=\sum_{k=1}^{K}\frac{I_{k}({\bm{z}}_{i})}{\sum_{j}I_{k}({\bm{z}}_{j})}\qquad I_{\text{Merge-GausNorm}}({\bm{z}}_{i})=\sum_{k=1}^{K}\frac{I_{k}({\bm{z}}_{i})-\mu_{k}}{\sigma_{k}}$
However, as shown in Tab. 3, both methods still underperform compared to our voting-based strategy, reinforcing the limitations of relying on score magnitudes directly.
Limitations of rank-based aggregation.  Rank-based methods sidesteps the challenge of comparing raw influence scores by focusing on within-task ranking. Round Robin selects samples by cycling through each task and picking the highest-scoring remaining sample for that task, ensuring balanced coverage. MinRank (Eqn. 5) selects samples that have the best minimum rank across all tasks, prioritizing examples that perform exceptionally well in at least one task regardless of their performance in others. Although these methods ensure balanced coverage across tasks, they can overfit to outlier tasks. This is particularly evident with LLaVA-W Bench llava , which is an outlier in its influence ranking: both Round Robin and MinRank achieve relatively high scores on it (e.g. MinRank: 68.4), but this comes at the cost of lower performance on all other tasks (Tab. 3). This suggests that purely rank-based selection can trade off some overall efficacy on the mainstream tasks and hurt multi-task balance. In contrast, our Vote approach is more robust and avoids this by focusing on multi-task consensus rather than forcing equal representation, yielding better balance and higher overall performance (98.6% Rel.), highlighting the importance of identifying broadly influential examples rather than narrowly optimizing per-task rankings.
Figure 3: Pairwise overlap heatmap between specialists. The values show overlap percentages between benchmarks’ selected samples.